OpenStudy (anonymous):

A packing operator can assemble 3 cardboard boxes in 6 minutes. At this rate, how many boxes can she assemble in 30 minutes? A. 18 B. 15 C. 9 D. 21

OpenStudy (anonymous):

Ill give medal!

OpenStudy (opcode):

Have you learned about proportions? Example: \[\dfrac{3}{6}=\dfrac{x}{30}\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

How do ya do it man.

OpenStudy (anonymous):

B is the correct ans

OpenStudy (anonymous):

Katie bought 6 hotdogs for $1.89 each and 3 drinks for $1.10 each. Which equation can be used to find the total cost, C, of the hotdogs and drinks? A. C = (6 × 1.89) + (3 × 1.1) B. C = (6 × 3) + (1.89 × 1.1) C. C = (3 × 1.89) + (6 × 1.1) D. C = (6 + 1.89) + (3 + 1.1)

OpenStudy (anonymous):

Last one

OpenStudy (opcode):

@ksaichandra I was typing the explanation out -_-. Please do not just give out answers.

OpenStudy (anonymous):

I want to know how to do it

OpenStudy (anonymous):

not just the answer

OpenStudy (anonymous):

guys????

OpenStudy (anonymous):

@Opcode

OpenStudy (opcode):

Okay, well basically there are things called a ratio. Normally written as: 1:2 or 3:6, etc. We can use ratios to set up proportions for bigger values. Example: \[\dfrac{3}{6}=\dfrac{x}{30}\] The 3/6 (3:6) is from the question. "3 cardboard boxes in 6 minutes" Now you have to cross multiply.

OpenStudy (anonymous):

I need to know this question man.....im stuck badly.....

OpenStudy (anonymous):

Katie bought 6 hotdogs for $1.89 each and 3 drinks for $1.10 each. Which equation can be used to find the total cost, C, of the hotdogs and drinks? A. C = (6 × 1.89) + (3 × 1.1) B. C = (6 × 3) + (1.89 × 1.1) C. C = (3 × 1.89) + (6 × 1.1) D. C = (6 + 1.89) + (3 + 1.1)

OpenStudy (anonymous):

Yes, I understand ratios...I didnt know how to set that up though....

OpenStudy (opcode):

Oh okay :-). Katie bought (6 hotdogs for $1.89 each and 3 drinks for $1.10 each). Which equation can be used to find the total cost, C, of the hotdogs and drinks? The soft brackets indicate what is important in the question. 6 hotdogs for $1.89 each 3 drinks for $1.10 each C (COST) = 1.89(h) + 1.10(d) h = 6 (hotdogs) d = 3 (drinks)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

....

OpenStudy (opcode):

Yes, A is correct. I hope you understand though.

OpenStudy (anonymous):

Thx guys :D

OpenStudy (anonymous):

there i voted :D